The researcher from Q2 decided that they want to follow up with the 6 children one year after they have completed the intervention program to see how they are now coping with negative emotions. The same method of assessment was used to examine coping responses. The data for the one-year follow-up list is listed below (along with results from Q2). Before One-Year Person Follow-U After Treatment Child Treatment Total Y 297 n=6 k 3 N 18 G 65 17 10 14 4 4 12 SSbefore 11.33

A. Use a repeated measures ANOVA with a=0.05 to determine whether there are any significant mean differences among the three conditions.

B. Calculate n2 to determine the effect size.

***PLZ SOLVE LEGIBLY AND WITH STEPS**
Answer #1


142 252 262 (65)2 18 = 249.50-234.72 -14.78 The sum of square totalis. (GT) SST -Sumof squareGT - 2 = 297-(45) = 297-234 .72The degree of freedom for between groups is, =3-1 The degree of freedom error is =18-3 -15 The degree of freedom total is, d,